d) Consider a CPU running a single process having an execution time of 25ns, in which 12 ns are for I/O process. Determine the CPU utilization.

(e) For the CPU in part (d), if the CPU is running a multiprogramming environment, and the CPU utilization needs to be doubled, find the number of processes that need to be running simultaneously.
Answer #1

A simple way to calculate CPU utilization is
U=R/C

U= Utilization
R= Requirements which in simple terms is the BUSY TIME
C= Capacity which is simple terms is BUSY TIME + IDLE TIME

For eg:
BUSY TIME= R = 3042450
BUSY TIME + IDLE TIME = C = 3042450 + 2644830 = 5687280

Here your ans will be = 25/(25+12)=67.5%

And second ans is 2 minimum two process need to be run